OpenStudy (anonymous):

Solve by using the perfect squares method. (3x + 3)2 = 36

OpenStudy (anonymous):

is the two supposed to be a square of the stuff in parentheses??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

would it be 12x^3

OpenStudy (gorv):

(3x+3)^2=36 take 3 common 3^2 * (x+1)^2=36 9* (x+1)^2=36 (x+1)^2=36/9 (x+1)^2=4 x+1=root 4 x+1= + -2 x+1=2 or x+1=-2 x=1 or x=-3

OpenStudy (gorv):

@Ashbtr18